# Diffusion Model in Latent Space for Medical Image Segmentation Task

### v2

@@ III-C Segmentation in latent spcce
Figure 2: The latent diffusion reverse process, where multiple segmentation masks are generated for a single input medical image.
@@ IV Experiments
IV-A Dataset
We conducted experiments on three datasets. The first dataset, ISIC-2018 [15], was used for the ISIC 2018 Task 1 challenge, which aimed to develop automated binary segmentation of skin lesion areas in dermatoscopic images. The ISIC-2018 dataset was obtained from the official competition website and was collected using web crawling techniques. The second dataset, CVC-Clinic [16], consists of multiple frames extracted from endoscopy videos and is commonly used for medical image segmentation, particularly for polyp detection. This dataset is publicly available on Kaggle, and we accessed it from this platform. The third dataset, LIDC-IDRI [17], contains lesion annotations from experienced thoracic radiologists. Unlike the first two datasets, LIDC-IDRI required specific preprocessing steps. The data was split patient-wise into training, validation, and testing sets in an 8:1:1 ratio. For the LIDC-IDRI dataset, slices containing nodules were extracted from patient cases. The four masks annotated by radiologists were included to evaluate the inter-rater uncertainty between the model’s predictions and the expert annotations. For the ISIC-2018 and CVC-Clinic datasets, the default data splits provided by the datasets were used.
@@ IV-B Implementation Details
We conducted experiments using VQ-VAE for both medical image and mask compression, and a diffusion model (DM) for segmentation in the latent space. The VQ-VAE employs an autoencoder architecture with three downsampling and upsampling layers, while the DM uses a UNet architecture with the same three-level downsampling and upsampling configuration. Both models use 128 base channels. For segmentation mask compression, a Weighted Cross-Entropy (WCE) loss is applied, with a weight of 5 for the positive class. For the LIDC-IDRI dataset, a higher weight of 50 is used due to the small size of the masks. The models are trained for 1000 steps using a simple loss function, following the settings in [5]. All experiments are implemented in PyTorch and conducted on four Tesla V100 GPUs, each with 16 GB of memory. The ISIC-2018 and CVC-Clinic datasets are uniformly resized to 256×256 pixels, while the LIDC-IDRI dataset is resized to 128×128 pixels. The networks are trained end-to-end using the AdamW optimizer [21]. The batch size is set to 32 for the ISIC-2018 and CVC-Clinic datasets and 64 for LIDC-IDRI. The initial learning rate is $1\times 10^{-4}$. During inference, all models utilize 5-fold ensemble predictions to enhance performance. During inference, all models utilize a 5-fold ensemble prediction approach to enhance performance. For a fair comparison, the diffusion-based competitor was also reproduced using the same settings. Additionally, commonly used evaluation metrics, including the Dice coefficient and Intersection over Union (IoU), are employed to quantitatively assess segmentation performance.
@@ IV-C Results
Figure 3: Illustration of our pipeline: two initial mask samples, their averaged output, and the confidence map.
@@
We conducted experiments sequentially for two tasks: (1) training a pair of VQ-VAE models for both medical images and masks to map them into the latent space, and (2) training a conditional diffusion-based model to learn the conditional latent mask distribution in this space. For the first task of perceptual data compression, VQ-VAE demonstrates strong capability in encoding, representing, and reconstructing the data. All results for the VQ-VAE are reported in Table I. For images, we used the default reconstruction loss (MSE). The evaluation metrics were consistently high across all three datasets, achieving an SSIM of 0.94 and PSNR of 34 on both ISIC-2018 and CVC-Clinic. The LIDC-IDRI dataset is more challenging, yielding only 0.89 SSIM and 33.2 PSNR. These results confirm the feasibility of mapping medical images into the latent space using VQ-VAE. For masks, the metrics were extremely high across all four evaluation measures, even when using only MSE loss. Dice and IoU reach 98.4 and 96 on ISIC-2018, while CVC-Clinic performs even better, with Dice and IoU exceeding 99. Although learning binary masks is generally an easier task, LIDC-IDRI produces slightly lower scores because its tiny masks can cause the model to overlook small structures. When switching to Weighted Cross-Entropy (WCE), performance improves across all three datasets. The gains on ISIC-2018 and CVC-Clinic are modest, but LIDC-IDRI benefits substantially, with Dice increasing from 88 to 94.4, IoU from 83.1 to 89.4, and SSIM and PSNR rising by 0.07 and 6.5, respectively. These results demonstrate the adjusted model’s capability to effectively map mask data into the latent space.
Figure 4: Dice results as the number of sampling steps increase.
For the second task, we provide qualitative results in Fig. 3 across the three datasets. We select five samples and compute their average to generate a confidence map, while using a threshold of 0.5 to visualize the ensemble mask consistency for all cases. For the quantitative analysis, we first examine how the evaluation metrics change as we increase the number of sampling steps, which can be interpreted as increasing the number of experts that the model imitates in a clinical setting. The Dice results are shown in Fig. 4. From 1 to 5 samples, the Dice score increases steadily, reflecting clear gains when generating more candidate masks. Specifically, ISIC-2018 improves from 0.865 to 0.880, CVC-Clinic from 0.81 to 0.845, and LIDC-IDRI from 0.834 to higher values. This trend is expected, as aggregating predictions from more experts typically produces better consensus. Beyond 5 samples, further increasing the number up to 30 yields only marginal benefits, with the curve becoming stable. Based on this observation, we choose 5 samples as the optimal setting, achieving a balance between computational efficiency and segmentation accuracy.
@@
Furthermore, we compared our model against baseline methods, covering both the traditional one-to-one approach and the newer one-to-many paradigm. First, Tab. II presents the results of models following the former approach. Common baselines such as UNet and UNet++ achieved relatively modest performance, whereas ResUNet and nnUNet demonstrated stronger capabilities, reaching around 87 Dice and 79 IoU on ISIC-2018, 82 Dice and 72 IoU on CVC-Clinic, and 82 Dice and 81 IoU on LIDC-IDRI. In contrast, our proposed model proves to be the most prominent, achieving superior Dice scores of 88, 84.5, and 83.4 (improvements ranging from 0.5 to 2.0) and IoU scores of 80.5, 73.1, and 71.8 (gains of approximately 0.8 to 1.8) across the three datasets. This highlights the effectiveness of our method over traditional one-to-one segmentation models.
@@
Compared with recent approaches, as shown in Tab. III, our proposed model outperforms all three baselines. EnsembleDiff and SegDiff show less stable performance than MedSegLatDiff. On ISIC-2018, MedSegDiff is slightly competitive with our method, achieving a Dice of 88.1 compared with our 88.0 and the same IoU of 80.5. However, on the other two datasets, our model clearly surpasses all baselines, reaching the best Dice of 84.5 (+0.5) and IoU of 73.1 (+0.6) on CVC-Clinic. A similar trend is observed on the more challenging LIDC-IDRI dataset with tiny masks, where our model achieves a margin of +0.9 Dice and +0.7 IoU over MedSegDiff. Overall, the findings highlight the clear superiority of our proposed model, which achieves the most robust and accurate segmentation results across all benchmarks.
@@ V Conclusion
In this paper, we proposed MedSegLatDiff, a conditional diffusion model operating in the latent space via VQ-VAE for medical image segmentation. We designed two VQ-VAE modules for images and masks to encode data into a compact latent representation. For perceptual mask compression, incorporating weighted cross-entropy as the reconstruction loss significantly improved the encoding–decoding quality. Our experiments show that using five samples yields stable and reliable results, enabling MedSegLatDiff to outperform previous one-to-one segmentation methods and compete effectively with recent one-to-many approaches. This strategy provides an efficient solution that improves both training and sampling efficiency, thereby enhancing segmentation performance through a sampling process that mimics a group of experts. For future work, we plan to explore more advanced conditioning mechanisms, such as classifier guidance [4], classifier–free guidance [22], and flow–based generative modeling, to further improve stability, calibration, and uncertainty estimation of the model’s outputs.

### v3

@@ IV Experiments
IV-A Datasets
We conducted experiments on three publicly available datasets. The first dataset, ISIC-2018 [39], was released for the ISIC Skin Image Analysis Workshop and Challenge 2018 (Task 1), which focuses on automated binary segmentation of skin lesion regions in dermoscopic images. The dataset was obtained from the official ISIC Challenge website [^1] https://challenge.isic-archive.com/data/\#2018 and consists of dermoscopic images collected from multiple clinical centers. The second dataset, CVC-Clinic [40], comprises frames extracted from endoscopy videos and is widely used for polyp segmentation in medical images. It is publicly available on Kaggle [^2] https://www.kaggle.com/datasets/balraj98/cvcclinicdb. The third dataset, LIDC-IDRI [41], consists of thoracic CT scans with lesion annotations provided by four experienced radiologists. Only slices containing nodules were extracted, and the four radiologist-annotated masks were averaged to form a single consensus mask used as the ground-truth target. Unlike the first two datasets, LIDC-IDRI was split on a patient-wise basis into training, validation, and testing sets with an 8:1:1 ratio. For ISIC-2018 and CVC-Clinic, the default dataset-provided splits were used.
@@ IV-B Implementation Details
We employ a VQ-VAE to compress both medical images and segmentation masks, followed by a diffusion-based model operating in the latent space for segmentation. The VQ-VAE adopts an autoencoder architecture with three downsampling and upsampling levels, while the diffusion model is implemented as a U-Net with a matching three-level configuration. Both networks use 64 base channels with channel multipliers $[1,2,4]$. For mask compression, a Weighted Cross-Entropy (WCE) loss is applied, with a positive-class weight of 5 for ISIC-2018 and CVC-Clinic, and 50 for LIDC-IDRI to account for the extremely small target regions. The diffusion model follows the training setup described in [22], using 1000 diffusion timesteps. To condition the diffusion process, the encoded input image is directly concatenated with the noisy latent representation, serving as guidance during both training and inference. Images from ISIC-2018 and CVC-Clinic are resized to $256\times 256$ pixels, while LIDC-IDRI slices are resized to $128\times 128$. Training is performed using the AdamW optimizer [42], with batch sizes of 32 for ISIC-2018 and CVC-Clinic and 64 for LIDC-IDRI. The initial learning rate is set to $1\times 10^{-4}$. During inference, five stochastic samples are generated per input and aggregated by averaging to obtain a confidence map, which is thresholded at 0.5 to produce the final segmentation. For fair comparison, diffusion-based baselines are reproduced under identical settings. Performance is evaluated using PSNR [45] and SSIM [44] for the reconstruction task, and the Dice coefficient and Intersection over Union (IoU) for the segmentation task. All experiments are implemented in PyTorch and conducted on four Tesla V100 GPUs with 16 GB memory each.
@@ IV-C Results
We first trained two separate VQ-VAE models for images and segmentation masks to enable latent-space mapping, and then trained a conditional diffusion model on the learned latent distributions. Table I reports the reconstruction performance of the VQ-VAE models
@@
For image reconstruction, we employed the mean squared error (MSE) loss, achieving high reconstruction quality with SSIM of 0.94 and PSNR of approximately 34 on the ISIC-2018 and CVC-Clinic datasets, while lower performance was observed on the more challenging LIDC-IDRI dataset (SSIM 0.89, PSNR 33.2). For mask compression, using MSE loss resulted in strong segmentation fidelity, with Dice scores exceeding 98% and IoU scores above 96% on ISIC-2018 and CVC-Clinic. However, the presence of small and fine-grained structures in LIDC-IDRI led to reduced performance, with a Dice score of 88.0%. When switching to the WCE loss, performance improvements on the ISIC-2018 and CVC-Clinic datasets are marginal, as their segmentation masks typically cover relatively large regions. In contrast, the LIDC-IDRI dataset shows substantial improvements, with Dice increasing by 6.4 points to 94.4, IoU improving by 6.3 points to 89.4, SSIM increasing by 0.07, and PSNR improving by 6.5 dB. Overall, only modest gains are observed on datasets with larger mask regions, while WCE proves particularly beneficial for datasets containing small and sparse structures. These results validate VQ-VAE’s effectiveness for latent-space compression of medical data, especially with WCE for sparse/tiny masks.
Figure 3: Dice scores versus the number of samples (ensemble size)
For the segmentation task, qualitative results of our model are first presented in Fig. 4. Five stochastic samples are generated for each input image and averaged to produce a confidence map, which is then thresholded at 0.5 to obtain the final segmentation mask. Results across all three datasets demonstrate that the generated masks closely match the ground-truth annotations, including the small target regions in the LIDC-IDRI dataset. Variations among sampled masks are mainly observed along object boundaries, which can be attributed to inherent ambiguities in medical images. By generating multiple segmentation masks, the proposed model produces a confident consensus prediction while explicitly capturing uncertainty through the confidence map. This uncertainty-aware output enhances the interpretability of the model and provides clinicians with additional insights for more informed analysis. Furthermore, we present the Dice score as a function of the number of samples in Fig. 3 to justify our choice of using five samples during inference. The segmentation performance improves steadily as the number of samples increases from 1 to 5, with Dice scores on ISIC-2018 rising from 0.865 to 0.880, on CVC-Clinic from 0.810 to 0.845, and on LIDC-IDRI from lower values to 0.834, indicating improved consensus as more stochastic samples are aggregated. Beyond five samples, performance gains gradually plateau (up to 30 samples), while computational cost continues to increase. Therefore, we adopt five samples as an effective trade-off between segmentation accuracy and inference efficiency.
Figure 4: Visualization of multiple segmentation masks. The first two columns show the input image and the ground-truth mask, followed by two mask samples, their averaged output, and the resulting confidence map.
@@
Moreover, quantitative comparisons between our method and existing approaches are reported in Tables II and III. Table II shows that MedSegLatDiff achieves competitive performance compared with one-to-one baselines. In particular, MedSegLatDiff obtains the highest Dice scores (88.0 / 84.5 / 83.4) and IoU (80.5 / 73.1 / 71.8) across all datasets, with consistent improvements ranging from 0.5 to 2.2 in Dice and from 0.5 to 0.8 in IoU compared to the best-performing competing methods. These results indicate the advantage of the proposed one-to-many modeling paradigm over conventional one-to-one approaches, highlighting its suitability for medical image segmentation tasks where ambiguity and inter-observer variability are inherent. In parallel, comparisons with one-to-many diffusion-based baselines are reported in Table III. MedSegLatDiff demonstrates competitive performance against other diffusion-based methods. Specifically, while it achieves comparable IoU and slightly lower Dice than MedSegDiff on ISIC-2018, it clearly excels on CVC-Clinic, improving Dice by 0.5 and IoU by 0.6. On the more challenging LIDC-IDRI dataset, where tiny and sparse masks pose greater difficulties, MedSegLatDiff further improves Dice by 0.9 and IoU by 0.7, highlighting its robustness in handling small target structures. Overall, these results demonstrate the effectiveness of latent-space diffusion with VQ-VAE compression and WCE in achieving robust and uncertainty-aware medical image segmentation.
@@
Limitation: Despite the promising results, our method still has several limitations. In the current study, we only integrate VQ-VAEs to enable latent-space segmentation and have not yet systematically investigated the invariance properties of different VAE variants for comparison. In addition, although our experiments demonstrate that WCE is more effective than MSE for mask compression, further exploration of loss functions specifically designed for tiny or sparse object segmentation, such as Focal Loss or Dice-based losses, is necessary. These aspects constitute important directions for future work and will be explored in our subsequent studies to further improve segmentation robustness and uncertainty modeling.
@@ V Conclusion
In this paper, we propose MedSegLatDiff, a conditional diffusion framework that operates in a compressed latent space using VQ-VAE for efficient medical image segmentation. By employing two separate VQ-VAE modules for images and segmentation masks, our method achieves effective perceptual compression while preserving critical structural details. Moreover, replacing the conventional MSE loss with a Weighted Cross-Entropy (WCE) loss for mask reconstruction substantially improves reconstruction quality, particularly for tiny and sparse structures such as nodules in the LIDC-IDRI dataset.
Our experiments demonstrate that generating five stochastic samples strikes a robust balance between computational cost and segmentation performance, producing stable ensemble predictions and reliable confidence maps. MedSegLatDiff consistently outperforms traditional one-to-one segmentation models and achieves competitive or superior performance compared to recent one-to-many diffusion-based approaches across ISIC-2018, CVC-Clinic, and LIDC-IDRI. The latent-space design effectively suppresses noise and allows the diffusion model to focus on the segmentation process, while the one-to-many paradigm captures segmentation uncertainty by mimicking inter-expert variability, thereby providing valuable support for clinical decision-making.
For future work, we plan to investigate advanced conditioning strategies, such as classifier guidance [25], classifier-free guidance [26], and flow-based generative modeling, to further enhance output stability, calibration, and uncertainty quantification.